OpenStudy (anonymous):

Finding the slope through this line. (-5,-3) and parallel to y=1

OpenStudy (anonymous):

(-5,-3) and (0,1) y-y0/x-x0 ==> 1-(-3)/0-(-5) = 4/5

OpenStudy (anonymous):

\[-5\left| 2x+3y \right|\] when x = 0.5 and y =-2 Evaluate

OpenStudy (anonymous):

Solve \[\sqrt{x+2} =\sqrt{x ^{2}}\] check for extraneous solutions.